OpenStudy (anonymous):

which expression gives the solutions of -5+2x^2=-6x ?

OpenStudy (anonymous):

((-b+/- sqrt(b^2 - 4*a*c))/2*a) this is the formula. 2x^2 + 6x - 5, in this, its basically ax^2 + bx - c Substitute :)

OpenStudy (anonymous):

bring over -6x \[2x^2+6x-5=0\] Now you can use quadratic equation to plug in the values and solve the equation. \[\frac{ (-b \pm \sqrt{b^2-4ac}) }{ 2a }\]

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

Youre welcome :)